### v2

# Pointing the Way: Refining Radar-Lidar Localization Using Learned ICP Weights

Abstract
This paper presents a novel deep-learning-based approach to improve localizing radar measurements against lidar maps. Although the state of the art for localization is matching lidar data to lidar maps, radar has been considered as a promising alternative. This is largely due to radar being more resilient against adverse weather such as precipitation and heavy fog. To make use of existing high-quality lidar maps, while maintaining performance in adverse weather, it is of interest to match radar data to lidar maps. However, owing in part to the unique artefacts present in radar measurements, radar-lidar localization has struggled to achieve comparable performance to lidar-lidar systems, preventing it from being viable for autonomous driving. This work builds on an ICP-based radar-lidar localization system by including a learned preprocessing step that weights radar points based on high-level scan information. Combining a proven analytical approach with a learned weight reduces localization errors in radar-lidar ICP results run on real-world autonomous driving data by up to $54.94\%$ in translation and $68.39\%$ in rotation, while maintaining interpretability and robustness.
I INTRODUCTION
Recent autonomous vehicle progress has been facilitated in part by the collection of detailed maps of commonly visited locations. When preconstructed maps are available, the primary vehicle navigation task is simplified to finding the vehicle pose within the map, commonly referred to as localization. An accurate and reliable localization estimate is critical for safe autonomous operation.
The current state of the art for the task of localization is matching live lidar scans to preconstructed lidar maps [1]. Lidar sensors are capable of capturing high-definition maps, which can then be used to achieve accurate localization. However, lidar measurements can be affected or completely blocked by precipitation and other challenging environmental conditions [2, 3, 4, 5]. Radar has been considered as an alternative for localization in all-weather conditions. Radar, owing to its longer wavelength, is much less susceptible to small particles in the air, and is thus largely unaffected by rain, snow, dust, or fog [6]. However, the longer wavelengths and lower resolution inherently lead to sparser maps as compared to those constructed using lidar. Additionally, radar measurements suffer from artefacts such as speckle (noisy measurements), saturation (falsely strong return signal), and ghosting (multiple echoes of the same point or object) [7]. These artefacts are challenging to model and heuristically extract [8]. As a result, radar-radar localization accuracy is still worse than that of lidar-lidar [2], making it not quite viable for tasks that require centimeter-level accuracy, such as autonomous driving.
Figure 1: A satellite image of the environment, overlaid with the full extracted radar pointcloud (left) and the radar pointcloud weighted using our method (right). The weight associated with each point in the weighted pointcloud corresponds to the intensity in colour. It can be seen that our method picks out some of the most important geometric features while ignoring noisy and unreliable points, thus facilitating the task of ICP.
Owing to the unique advantages and disadvantages of lidar and radar, there has been interest in multi-modal radar-lidar localization. By localizing using radar scans against already available high-quality lidar maps, the goal is to achieve near state-of-the-art localization performance in all weather conditions. However, past work in radar-lidar localization has not been able to achieve the accuracy required for autonomous driving [2, 9, 5, 10]. Our previous work, [2], was able to achieve state-of-the-art radar-lidar localization by using the iterative closest point (ICP) algorithm to match radar and lidar pointclouds, but still fell slightly short of desired localization accuracy [11].
Our hypothesis is that unique radar artefacts and noise properties significantly degrade the quality of radar-lidar localization. To combat this, we refine ICP-based radar-lidar localization using a learned weight mask to identify only the most useful radar points for the task of radar-lidar localization. This learned filtering approach, visualized in Figure 1, is the primary contribution of this paper. Additionally, we provide an open-source, stand-alone, differentiable ICP library for future approaches that wish to incorporate learned elements in an ICP pipeline.
The rest of this paper is structured as follows. Section II presents relevant prior work. Section III details our approach. Finally, Section IV provides experimental results, with concluding remarks presented in Section V.
II RELATED WORK
Radar has seen a resurgence within robotics in recent years [12, 13, 14]. The ability of modern-day radar systems to produce reasonable long-term odometry results is shown in [15], where an ICP-style pointcloud alignment is used to compute the relative motion between consecutive timestamps. An alternative approach is presented in [16], where a network is trained to mask raw radar scans before using them for correlative scan matching of consecutive frames. It is shown that such a learned approach is capable of outperforming heuristic filtering methods. A network that directly extracts features, descriptors, and feature scores from raw radar images and then uses them to match radar data from consecutive timesteps is presented in [17]. In our previous work [18], features are learned in a self-supervised manner, without needing groundtruth data. Although feature-extraction methods are typically more interpretable and robust to poor initial guesses, they frequently require large networks with higher computational cost. Additionally, feature matching across sensing modalities, such as between radar and lidar, poses an extra challenge. Finally, it is important to note that most radar algorithms have focused on odometry or simultaneous localization and mapping (SLAM), e.g. [19, 20, 21, 22], meaning that investigations into the performance of radar in localization against a previously known map is understudied. An extended Kalman filter (EKF) ICP-based approach for radar-radar localization is presented in [23]. However, their localization errors are on the order of meters, far above autonomous driving requirements. In our recent ICP-based approach [2], localization is done using batch optimization with a modified Constant False Alarm Rate (CFAR) [24] point extractor called Bounded False Alarm Rate (BFAR) [25]. This work shows that radar localization against an existing radar map can achieve promising results that approach the level needed for autonomous driving.
Radar has also been used to localize against lidar maps. One of the first algorithms for radar-lidar localization was proposed for indoor disaster environments where lidar is blocked by smoke [5]. This algorithm implements a recursive undistortion approach to utilize shape matching between live radar scans and an existing lidar map. This recursive implementation is shown to overcome some aspects of radar noise and artefacts to achieve usable localization results in situations where lidar measurements completely fail. A learning-based approach was taken in [26], where a generative adversarial network (GAN) was used to style-transfer a live radar scan to lidar data. Instead of doing a cross-sensor style transfer, Yin et al. [9] use a set of networks to transfer both radar and lidar scans to a common learned embedding, in which they then perform cross-correlation matching. This match is then treated as a measurement in a Kalman-filter framework. Their algorithm is set up to be fully differentiable, meaning that the networks can be trained to produce an embedding that directly minimizes the final localization output from the Kalman filter. Another embedding-style approach is presented in [10], where it is hypothesized that although extracted radar and lidar points do not necessarily have one-to-one correspondences, their groupings are related. Thus, radar points are binned into discrete Cartesian and polar groupings and an initial alignment between radar and lidar groupings is done via a density metric. The estimate is then further refined using ICP. A learned radar-lidar place recognition algorithm is presented in [27], but they do not evaluate localization accuracy. Although promising, none of these papers are able to achieve centimeter-level localization accuracy, a requirement for safe autonomous vehicle operations. However, our previous work was able to achieve an accuracy on the order of tens of centimeters with a nonlearned approach [2]. This algorithm first converts the radar and lidar measurements to pointclouds using BFAR, and then applies ICP to match the pointclouds.
In our previous work, we have found that CFAR-based point detectors work well for generating radar pointclouds. These can subsequently be matched to prebuilt lidar maps using ICP to yield state-of-the-art radar-lidar performance. This approach makes use of the full-resolution radar scan without needing to downsample, as in [18], and does not require any discretization of the pose estimate, as in [9, 16]. However, CFAR-like detectors only use local information to extract points, meaning that a significant amount of noise gets extracted. Thus, our approach is to start with the radar-lidar localization pipeline presented in [2] and, building on [16], incorporate a learned weight mask into the ICP algorithm to identify only the most useful radar points for the task of radar-lidar localization. Using a network allows us to learn the high-level spatial contextual cues in order to softly reject points that may correspond to noise, while still retaining the high-resolution, discretization-free point localization from CFAR and ICP.
III METHODOLOGY
This section presents the various parts of the pipeline. First, the overall training architecture is presented. Then, the differentiable ICP (dICP) and supervisory map mask components are discussed in more detail. Our code base is publicly available to facilitate reproducibility and extensions [^1] Paper code is available at: https://github.com/lisusdaniil/mm_masking..
Figure 2: The pipeline used to train the weight mask network. A radar scan, used to localize against a reference lidar map, is fed to a U-Net network that produces a weight mask. In parallel, the scan is used to extract a pointcloud using the BFAR detector. The pointcloud then indexes weights from the weight mask. The points and corresponding weights are used in a weighted differentiable ICP algorithm (dICP) to localize against the reference map pointcloud. The training loss is formed through a term corresponding to the error in the dICP result and a binary cross-entropy (BCE) term computed between the weight mask and a supervisory map mask. The network is trained by backpropagating the loss through dICP and the weight extraction.
III-A Architecture
The overall architecture is presented in Figure 2. The input to the pipeline is a single radar intensity scan. A Cartesian representation of the scan, normalized to have a maximum pixel value of one, is used as input to the network, whereas an unnormalized polar representation is fed to the point-extraction module. Inspired by [16], we use a U-Net network [28] to learn the weight mask for each radar input. A U-Net lends itself well to our application, as we are, in essence, trying to segment the radar scan into ‘good’ and ‘bad’ areas for pointcloud extraction. The encoder increases the number of channels from one, the measured intensity, to 8, 16, 32, 64, 128, and 256 channels. Each channel increase is composed of convolution blocks that include a 2D convolutional layer (kernel size 3, padding 1), a ReLU activation, another 2D convolution, a dropout layer with probability of 0.05, and finally a max-pooling operation (kernel size 2, stride 2). Using dropout helps to prevent over-fitting and encourages the network to learn higher-level, general information instead of relying on specific pixels [29]. The decoder then decreases the channel number from 256 down to 8 using the same steps as the encoder. For each decrease, the input is up-sampled, passed through the same convolution block as described for the encoder, concatenated with the corresponding skip connection from the encoder section, and passed through a convolution block again. Finally, the output is passed through one last 2D convolutional layer to decrease the 8 channels back down to one, and then through a sigmoid activation. The resulting mask is then normalized to have a maximum value of one. The network is implemented in PyTorch [30].
In parallel to the weight-mask generation, a pointcloud is extracted from the unnormalized, polar radar intensity scan using the BFAR detector. This pointcloud is then used to index the weight mask using bilinear interpolation between pixels. The extracted weight values, alongside the radar pointcloud, are then fed into the weighted dICP algorithm to be aligned with a given lidar map pointcloud.
Note, that as radar is a 2D sensor, the overall pipeline runs in 2D. This means that the dICP algorithm estimates a transform from the radar scan to the lidar map, ${\hat{\mbf{T}}}_{lr}\in SE(2)$, which is contrasted with the groundtruth pose, $\mbf{T}_{lr}\in SE(2)$, to form the error vector $\mbf{e}$ according to
$\mbf{e}=\begin{bmatrix}{e}_{x}&{e}_{y}&{e}_{\phi}\end{bmatrix}^{\mathsf{T}}=\log\left({\hat{\mbf{T}}}_{lr}\mbf{T}_{lr}^{-1}\right)^{\vee},$ (2)
where $\log(\cdot)$ is the logarithmic operator that converts $SE(2)$ elements to their Lie algebra representation and $(\cdot)^{\vee}$ is the operator that converts the Lie algebra representation to $\mathbb{R}^{3}$ [31]. Additionally, ${e}_{x}$, ${e}_{y}$, and ${e}_{\phi}$ are the $x$ (longitudinal), $y$ (lateral), and $\phi$ (heading) error components, respectively.
The loss is then formed based on two signals. The first, $\mathcal{L}_{\mathrm{ICP}}$, is based on $\mbf{e}$, and is formed as
$\mathcal{L}_{\mathrm{ICP}}=\mbf{e}^{\mathsf{T}}\begin{bmatrix}\alpha&0&0\\ 0&\alpha&0\\ 0&0&\beta\end{bmatrix}\mbf{e},$ (6)
where $\alpha,\beta\in\mathbb{R}$ control the relative impact of each pose term.
The second loss signal, $\mathcal{L}_{\mathrm{BCE}}$, is a binary cross-entropy loss computed between the generated weight mask and a supervisory map mask discussed in Section III-C. The final loss term is then of the form
$\mathcal{L}=\mathcal{L}_{\mathrm{ICP}}+\gamma\mathcal{L}_{\mathrm{BCE}},$ (7)
where $\gamma\in\mathbb{R}$ controls the relative impact of $\mathcal{L}_{\mathrm{BCE}}$ on $\mathcal{L}$.
After training is complete, weight masks can be generated from single radar scans during inference time. The subsequent weighted pointcloud can be fed to any ICP algorithm.
III-B Differentiable ICP
A key part of the pipeline is the dICP algorithm. This algorithm is made publicly available222dICP code is available at: https://github.com/lisusdaniil/dICP, and adds to the recent corpus of open-source, differentiable versions of classic robotic algorithms [32, 33, 34]. Indeed, all three of these libraries include a differentiable ICP implementation. However, they are all restrictive in that they only implement either point-to-point or point-to-plane cost terms, do not make use of any or only allow for one type of robust cost function, do not accept weighted pointclouds, and, for [33, 34], are embedded within a SLAM pipeline, without a clear way to run ICP in a stand-alone fashion.
Our dICP implementation, based in PyTorch, is entirely stand-alone and light-weight, making it simple to plug into any project requiring a differentiable ICP component. We provide support for both point-to-point and point-to-plane ICP. We also include options for either a Huber or Cauchy loss and make it easy to add additional losses. It is also possible to specify a trimming parameter to run trimmed ICP, and to provide weights for the source pointcloud in order to run weighted ICP. Depending on the problem at hand, it is possible to toggle dICP to run in 2D or 3D, where 2D mode prevents the algorithm from aligning the pointclouds in roll, pitch, and $z$ degrees of freedom. Finally, the differentiability of the algorithm can be toggled on and off, in case it is desired to run approximation-free ICP during inference. The main components of dICP are described in the following subsections.
III-B1 Nearest Neighbour
The first important step of each ICP iteration is to find the nearest point in the target pointcloud from each source pointcloud point. Here, ‘nearest’ refers to the smallest $L_{2}$ norm distance. Although computing the $L_{2}$ distance between each source point and every map point is differentiable, picking out the minimum index from this result is not. Of note is the fact that the act of indexing is differentiable if the index remains constant, as this amounts to a constant matrix multiplication. However, since the nearest neighbour for each point changes as ICP iterates, the index is not constant. One popular approximation for the $\mathrm{argmax}$ operation, or $\mathrm{argmin}$ in our case, is the Gumbel-Softmax [35, 36]. This approximation turns the distances into probabilities and then computes a softmax across these probabilities to find the ‘soft’ nearest neighbour. However, as this ‘soft’ neighbour is computed based on an average across all probabilities, it ceases to correspond to a true, physical map point, potentially limiting how finely the source pointcloud may be aligned to the target.
Instead, the dICP algorithm treats finding the nearest neighbour as a locally constant operation, and still makes use of $\mathrm{argmin}$. Consider one source point $\mbf{s}$ that is initially well aligned with the target point $\mbf{t}_{1}$ to which it actually corresponds. Shifting $\mbf{s}$ in the neighbourhood of $\mbf{t}_{1}$ will continue to yield $\mbf{t}_{1}$ as the nearest neighbour, meaning the operation can be viewed as constant. Moreover, gradient descent will be effective in minimizing the distance between $\mbf{s}$ and $\mbf{t}_{1}$. Now, consider that $\mbf{s}$ is instead in the vicinity of a different target point $\mbf{t}_{2}$. Alone, $\mbf{s}$ will locally converge to $\mbf{t}_{2}$ if gradient descent is run. However, we make the assumption that on average, the source pointcloud will be drawn in the correct direction towards the target through local convergences to neighbours. This means that other source points, locally converging to intermediate target points, will eventually drive $\mbf{s}$ into the neighbourhood of $\mbf{t}_{1}$. Finally, when the source points become aligned well enough, the nearest-neighbour operation becomes truly constant and a fine alignment becomes possible. Although this approach has the risk of leading to worse gradients and slower learning, it requires no tuning and remains identical between training and inference. Simple learning problems were used in order to validate this approach.
This approach performs well in the context of this paper. However, studying the tradeoff between using an approach like Gumbel-Softmax and treating the nearest-neighbour operation as constant is left for future work.
III-B2 Error Computation
Using the nearest neighbours, an error term based on either point-to-point or point-to-plane is computed. This operation is fully differentiable.
III-B3 Trim Distance
If a trim distance is specified, the difference between the norm of the distance between each point and its nearest neighbour is computed. This difference is then fed into a $\mathrm{tanh}$ function transformed to scale inputs between zero and one. The result is a weight smoothly transitioning from one for points that are within the trim distance to zero for points that are outside. A threshold is used when differentiability is turned off.
III-B4 Robust Loss
Next, a robust-loss function can be specified. A pseudo-Huber loss is used for the Huber loss during differentiable computations [37], while the Cauchy loss remains the same as it is already differentiable.
III-B5 Update Step
Finally, a weighted least-squares problem is formed using all error terms with weights corresponding to the multiplication of the trim weight with the robust-loss function weight. If initial weights are provided, such as those generated by our weight mask, then they are multiplied in as well. Gradient descent is used to update the estimated transform. If 2D ICP is required, extra dimensions in the update step are set to zero. Gradient descent is, by definition, differentiable and thus poses no issues.
III-B6 Iteration
Finally, the above steps are repeated until convergence or a maximum number of iterations is reached. Each iteration simply adds additional links in the PyTorch computational graph, meaning that any number of iterations are possible. However, in practice, the number of iterations may be restricted by available memory, as a computational graph is maintained with respect to all input points.
III-C Supervisory Map Mask
In order to better constrain the learning and help initial convergence, a supervisory signal for the weight mask based on the reference map is added to the loss term. To generate this reference, each map point is first transformed into the radar sensor frame using the groundtruth. All points in the range of the radar scan are then projected to a Cartesian grid, with pixel values nearest to the projected points being set to one, while all others are set to zero. This ‘map mask’ is then compared to the weight mask produced by the network through a binary cross-entropy loss, leading to the loss term $\mathcal{L}_{\mathrm{BCE}}$ in (7). This is found to help stabilize the learning process and ensure a more robust weight mask.
IV EXPERIMENTS
This section describes the dataset used for experiments, presents which architecture parameters were used, and discusses the results of applying our learned weight mask in the context of autonomous vehicle radar-lidar localization.
[…]
1 0.098 0.143 0.288 85.59 0.088 0.112 0.217 91.51 0.257 22.17 24.83
2 0.253 0.317 1.204 83.70 0.114 0.157 0.381 91.30 54.94 50.41 68.39
3 0.591 0.824 4.132 73.04 0.314 0.476 2.002 88.47 46.85 42.29 51.55
4 0.961 1.275 8.030 57.55 0.614 0.909 4.850 80.78 36.18 28.73 39.60
5 1.290 1.653 12.26 45.70 1.006 1.405 8.948 68.28 22.01 14.98 27.00
IV-A Dataset
The dataset used to train, validate, and test our approach is generated using the topometric localization pipeline from [2]. This pipeline is based on the VT&R framework [38], where a sequence is used to construct a series of connected submaps against which repeated traversals are localized.
This architecture is tested using the Boreas dataset [39], in which the same path is repeatedly traversed over the course of a year by a car equipped with radar and lidar sensors. To generate samples for this paper, the topometric localization pipeline was used to construct lidar submaps based on one traversal. During subsequent traversals, radar was used as the ‘live’ sensor. Whenever the pipeline ran ICP for radar-lidar localization, the radar scan, radar pointcloud, and map pointcloud were saved. Radar points were extracted using $a~{}=~{}1.0,b~{}=~{}0.09$ BFAR parameters, as per [2].
IV-B Training, Validation, and Testing
To mirror [2], we used point-to-point ICP with a trim distance of $5.0$\mathrm{m}$$ and a Cauchy robust loss with a parameter of $1.0$. We trainedu using the Adam [40] optimizer with a learning rate of $1e^{-5}$ and a batch size of 5. Approximately $50,000$ samples from 12 trajectories were used for training. The loss weighting parameters were set to $\alpha=\beta=\gamma=1$. Since we wish to improve on localization results that already achieve a low error, we initialized all initial guesses to the groundtruth. More precisely, we used the groundtruth to transform the map pointcloud into the radar frame, and then used identity as the initial guess for all samples. This additionally helped to keep the number of dICP iterations at 10 during training, greatly reducing the memory requirements and improving epoch speed. To encourage the network to learn orientation-agnostic masks, the radar scan and aligned pointclouds were randomly rotated between $0$\mathrm{\SIUnitSymbolDegree}$$ to $360$\mathrm{\SIUnitSymbolDegree}$$. This random augmentation happened every time a sample was loaded, meaning that the same localization pair was rotated differently every epoch.
Validation used $2,000$ samples taken from 3 additional trajectories. The differentiability in dICP was turned off to match practical ICP algorithms, but the initial guess was still initialized to the groundtruth. In order to ensure that ICP converged for validation, the maximum number of iterations was increased to 100.
Finally, the best weights were used to run tests on 6 full additional trajectories, testing approximately $27,000$ frames. dICP was run in nondifferentiable mode and given 100 iterations to converge. To study the effects of the learned weight mask in practical localization applications, the lateral, longitudinal, and heading components of the initial guess were sampled from uniform distributions with varying bounds. The bounds are ±$\pm 0.5\sigma$\mathrm{m}$$ for the translation components and ±$\pm 0.1\sigma$\mathrm{rad}$$ ($\approx 5.73$\mathrm{\SIUnitSymbolDegree}$$) for the heading, for noise scales $\sigma=\{1,\dots,5\}$. To evaluate the impact of using our learned weight masks, ICP was run twice: once with all point weights set to one as a baseline (i.e. the same as not using a mask), and once with weights sampled from weight masks generated by the network.
Figure 3: Two examples of the output from the pipeline, with one shown per row. Of note is the fact that the mask appears to be focusing on finding good clusters on which to run point-to-point ICP, as opposed to long geometric shapes such as walls. This is hypothesized to be due to the difficulty of point-to-point to match points along walls. A substantial amount of points without a map association, likely extracted from radar noise, are removed.
IV-C Results
Two sample masks and their qualitative effects are shown in Figure 3. Figure 4 shows the distribution of errors in each state component for the baseline and our method. Table I presents the root mean squared error (RMSE) results for each state component for the baseline and our method. It also presents the percent decrease in RMSE across all state components when using our method compared to the baseline. Finally, it shows the percent of ICP results that converged to within $0.2$\mathrm{m}$$ translation and $1$\mathrm{\SIUnitSymbolDegree}$$ rotation of the ground truth. These bounds were chosen as representative of an appropriate safety envelope for autonomous driving. As can be seen, our method improves the RMSE in every component at every noise scale. The effect is most pronounced for a noise scale of two, at which initial guesses were sampled from a uniform distribution with translational bounds of ±$\pm 1$\mathrm{m}$$ and heading bounds of ±$\pm 0.2$\mathrm{rad}$$ ($\approx 11.46$\mathrm{\SIUnitSymbolDegree}$$). At this noise scale, our learned weights improved the translation result by up to $54.94\%$ and rotation by $68.39\%$, maintaining the translation RMSE at approximately 10 centimeters. Our method also generally reduces the spread of errors in all components. This effect is more pronounced at higher scales, implying that our method improves robustness to poor initialization in addition to overall accuracy. The improvement in robustness is also shown in the increase in the percent of converged ICP results for all noise scales. One case we failed to meaningfully improve was the lateral component for a noise scale of one. This is likely due to the nature of the autonomous driving context: the presence of buildings and road-side objects constrains the lateral direction well even in the presence of noise, especially when a great initial guess is used.
Figure 4: Box plot showing the distribution of signed errors between the baseline and our method in lateral translation, longitudinal translation, and heading. The results are presented for a changing noise scale, which is a factor applied to the boundaries of a uniform distribution from which initial guesses are sampled. Our method has a tighter error distribution in all components and across different noise scales compared to the baseline, with the improvement more pronounced at higher initial errors.
Interestingly, the percent improvement of using our method first increases and then slowly drops in value towards higher noise scales. We hypothesize that the removal of points that are helpful for fine alignment is not necessarily beneficial for coarse alignment. Indeed, some points that are critical for coarse alignment may be removed in the process. However, our method continually improves the average accuracy in all components of radar-lidar localization across a variety of initial guesses.
Intuition for what the network is doing can be seen in Figure 3. As we run point-to-point ICP, long geometric features such as walls do not necessarily help constrain the translation. As a result, the network seems to learn to focus in on stand-alone features and corners, which make it easy to correctly associate nearest neighbours in point-to-point ICP. We are nonetheless able to see substantial improvement from both good and bad initial guesses, implying that extracting a smaller number of consistently good points is sufficient to improve the average ICP performance from a variety of initial conditions. A detailed study about how the masks change with point-to-plane ICP is left for future work.
V CONCLUSION
This paper provides a novel method for improving radar-lidar localization by learning weights associated with radar pointclouds to be used in ICP. This approach makes use of a localization pipeline with a heuristic point extractor and ICP. Instead of replacing any functionality with an end-to-end network, we use a network to generate a weight mask to be sampled by the heuristically extracted pointcloud. The sampled weight is then used to control the value of radar points in a weighted ICP. To train this network, we additionally introduce a publicly available novel implementation of a differentiable ICP algorithm. Our approach ensures that the robustness and interpretability of an analytical pipeline is maintained, while reducing error compared to using a purely heuristic point extractor up to $54.94\%$ in translation and $68.39\%$ in rotation in isolated radar-lidar localization tests. This makes radar-lidar localization more feasible to use in autonomous driving, providing greater safety and robustness in all weather conditions.
Future work will focus on improving this method through additional supervisory signals to prevent outliers that sometimes make the maximum error worse. Additionally, it is desired to integrate the learned weights into a full localization pipeline to see their impact in the presence of priors on ICP. Finally, it is of interest to evaluate the generalizability of this network to new maps, or to different radar sensors on the same map.
ACKNOWLEDGMENT
Resources used in preparing this research were provided, in part, by the Province of Ontario and the Government of Canada through their support of the Vector Institute.